OpenStudy (baby456):

fan and medal help now .Frieda is keeping track of how many coupons she has saved on a piece of paper. She spilled some coffee on the paper.her list now looks like this blank represents where the coffe spill occured.started :30 coupons, 1 week: coupons 2 weeks: after 3 weeks :240 coupons. a. assume that the sequence of coupons is arithmetic.how many coupons did she have after one week,2 weeks? b.assume that the sequence of coupons is geometric.how many coupons did she have after 1 week, 2 weeks?

OpenStudy (campbell_st):

ok... so the pattern is Week | start | 1 | 2 | 3 | 4 -------------------------------------- Coupons| 30 | ? | ? | 240 | (a) if the sequence is arithmetic then the same value is added each week this is called the common difference which can be found by d = (240 -30)/3 then d = 70 so in week 1 she has 30 + 70 = Week 2 she has 30 + 70 + 70 = see if you get the correct value for week 3 (b) for a geometric sequence you need \[a_{n} = a \times r^{n -1}\] you'll need to find r given \[a_{4} = 240... a = 30,... and..... n = 4\] the reason n = 4 is there there are 4 pieces of data start, then weeks 1, 2, 3 so find r \[240 = 30 \times r^{4 -1}\] \[8 = r^3\] so the common ratio is r = 2 so the pattern is start = 30 week 1 = 30 x 2 = week 2 = week 1 x 2 = week 3 = week 2 x 2 = see if it fits the pattern

OpenStudy (campbell_st):

I hope it all makes sense

OpenStudy (baby456):

thanks

OpenStudy (baby456):

can you help mew with another question

OpenStudy (campbell_st):

sorry I have to go and do some jobs

OpenStudy (baby456):

bye